# A Tutorial on Beyond-Diagonal Reconfigurable Intelligent Surfaces: Modeling, Architectures, System Design and Optimization, and Applications

### v1

@@ III-C.2 Fully-Connected
Figure 8: Examples of reciprocal architectures: (a) A 4-element D-RIS with single-connected architectures; (b) a 4-element BD-RIS with fully-connected architectures; (c) a 6-element BD-RIS with group-connected architectures (each group containing 3 interconnected elements); (d) a 4-element tridiagonal BD-RIS with tree-connected architectures; (e) a 5-element arrowhead BD-RIS with tree-connected architectures; (f) a 6-element BD-RIS with forest-connected architectures (each group containing 3 interconnected elements); (g) a 4-element BD-RIS with band-connected architectures (band width $q=2$); (h) a 5-element BD-RIS with stem-connected architectures (stem width $q=2$).
III-C.3 Group-Connected
The circuit complexity (characterized by the required number of reconfigurable admittance components) of the fully-connected architecture grows quadratically with the number of elements $M$. This makes it difficult to implement fully-connected BD-RISs in practice especially when a large dimension is required. To trade the circuit complexity and wave manipulation flexibility of RIS, a group-connected architecture has been proposed in [26]. In this architecture, the $M$ ports in the reconfigurable impedance network are uniformly divided into $G$ groups, and every $\bar{M}=\frac{M}{G}$ ports within one group are connected to each other to construct a fully-connected architecture. Specifically, in group $g$, $\forall g\in\mathcal{G}=\{1,\ldots,G\}$, each port $m_{g}=(g-1)\bar{M}+m$ is connected to ground via it own reconfigurable admittance component $Y_{m_{g}}$ and to another port $m^{\prime}_{g}=(g-1)\bar{M}+m^{\prime}$, $m^{\prime}\neq m$, $\forall m,m^{\prime}\in\bar{\mathcal{M}}=\{1,\ldots,\bar{M}\}$, via $Y_{m_{g},m^{\prime}_{g}}$, which satisfies $Y_{m_{g},m^{\prime}_{g}}=Y_{m^{\prime}_{g},m_{g}}$ for $m^{\prime}>m$. This indicates that the group-connected architecture is realized by in total $G\frac{\bar{M}(\bar{M}+1)}{2}=\frac{M(\bar{M}+1)}{2}$ reconfigurable admittance components. An illustrative example for a 6-element BD-RIS with group-connected architecture is given in Fig. 8(c). Hence, $\mathbf{Y}_{I}$ is a block-diagonal matrix
@@ III-C.3 Group-Connected
where each $\mathbf{Y}_{I,g}\in\mathbb{C}^{\bar{M}\times\bar{M}}$ is a full and symmetric matrix. Accordingly, $\mathbf{Z}_{I}$ is a block-diagonal matrix with each block being a full and symmetric matrix. By (17) and (19), for a lossless network, its scattering matrix $\mathbf{\Theta}$ is a block-diagonal matrix, with each block, $\mathbf{\Theta}_{g}\in\mathbb{C}^{\bar{M}\times\bar{M}}$, being symmetric and unitary, i.e.,
@@
III-C.4 Tree-Connected
While fully-connected architecture has the highest wave manipulation flexibility to achieve optimal performance, this is achieved at the cost of complicated circuit topology design requiring a large number of reconfigurable admittance components. To reduce the circuit complexity while maintaining the optimal (i.e., same as fully-connected) performance in single-user multiple input single output (MISO) systems, tree-connected architecture where the circuit topology forms a tree based on graph theory [39] has been proposed in [40]. In this architecture, the circuit complexity reduces to in total $12M-$ 1 admittance components, which can be much less than that of the fully-connected architecture. Two typical tree-connected architectures, leading to respectively tridiagonal and arrowhead admittance matrices, have been introduced in [40] and are reviewed as follows.
  • Tridiagonal: In this case, each port $m$ is connected to ground via its own admittance $Y_{m}$ and to port $m+1$ via an admittance $Y_{m,m+1}$, $\𝑀\forall m\in\mathcal{M}\backslash\{M\}$. An illustrative example for a 4-element BD-RIS with such architecture is given in Fig. 8(d). This mathematically leads to a tridiagonal admittance matrix
@@ III-C.4 Tree-Connected
III-C.5 Forest-Connected
Similar to the extension from fully-connected to group-connected architectures, the circuit complexity of the tree-connected architecture can be further reduced by dividing the $M$ ports into $G$ groups and constructing each group as a $\bar{M}$-port tree-connected architecture. This is referred to as the forest-connected architecture [40] with a circuit complexity $G(2\bar{M}-1)$. An illustrative example for a 6-element RIS with forest-connected architecture is given in Fig. 8(f). Hence, $\mathbf{Y}_{I}$ is a block-diagonal matrix with each block being a symmetric and tridiagonal (or arrowhead) matrix. Two extreme cases with $G=M$ and $G=1$ respectively correspond to single-connected and tree-connected architectures. For a lossless network with the forest-connected architecture, its impedance matrix $\mathbf{Z}_{I}$ also has a block-diagonal shape with each block being full and symmetric, and its scattering matrix $\mathbf{\Theta}$ is generally a block-diagonal matrix with each block being symmetric and unitary. Therefore, using $\mathbf{\Theta}$ or $\mathbf{Z}_{I}$ is, gain, not sufficient to fully reflect the mathematical constraint of the forest-connected architecture.
III-C.6 Band- and Stem-Connected
Inspired by the tree-connected architecture that can perfectly match the performance of fully-connected architecture in single-user MISO systems with significantly reduced circuit complexity, the optimal architectures that can perfectly match the performance of fully-connected architecture in multi-user MIMO systems have been proposed in [41]. In the optimal architectures, each port is connected to ground via its own reconfigurable admittance component, while the number of interconnections between ports is theoretically constrained by the DoF in multi-user MIMO systems, and is generally much less than that of fully-connected architectures. Two representative examples which could construct optimal architectures, namely band-connected [41] and stem-connected [41, 42], are reviewed as follows.
  • Band-Connected: In this case, each port $m$ is connected to ground via $Y_{m}$ and to the following $q$ ports, i.e., $m+1,\ldots,m+q$ (if any), via corresponding admittance components, where $q$ denotes the band width. An illustrative example for a 4-element BD-RIS with band-connected architecture ($q=2$) is given in Fig. 8(g). This mathematically leads to a band admittance matrix
@@ III-C.6 Band- and Stem-Connected
  • Stem-Connected: In this case, each port is connected to ground via $Y_{m}$. Meanwhile, there are $q$ ports with $q$ denoting the stem width, marked as $c_{i},\forall i\in\{1,\ldots,q\}$, each of which connects to all other ports via corresponding admittance components. An illustrative example for a 5-element BD-RIS with such architecture ($q=2$, $c_{1}=1$, $c_{2}=2$) is given in Fig. 8(h). Assuming $c_{i}\in\{1,\ldots,q\}$, we have an admittance matrix with the following structure:
@@
It is worth noting that, similar to tree- and forest-connected architectures, for a lossless network with band- and stem-connected architectures, its impedance matrix $\mathbf{Z}_{I}$ and scattering matrices $\mathbf{\Theta}$ will both be full matrices. Therefore, $\mathbf{Y}_{I}$ should be used to accurately reflect the constraint of these two architectures.
III-C.7 Dynamically Connected
@@
In addition to the aforementioned fixed architectures which do not adapt to the channel environment (i.e., the admittance components are tuned but the circuit topology dictated by the interconnections is fixed), a dynamically connected architecture whose circuit topology can be changed during the transmission has also been proposed in [43]. This can be implemented by an admittance-switch network which enables a joint control of the values of admittance components and the ON/OFF states of switches. In this architecture, every port $m$ is connected to ground via its own admittance $Y_{m}$ and connected to port $m^{\prime}$, $m^{\prime}>m$, $\forall m,m^{\prime}\in\mathcal{M}$ by the series of an admittance $Y_{m,m^{\prime}}$ and a switch $\mathsf{SW}_{m,m^{\prime}}$. This indicates that the dynamically connected architecture is realized by in total $\frac{M(M+1)}{2}$ reconfigurable admittance components and $\frac{M(M-1)}{2}$ switches. An illustrative example for a 6-element RIS with dynamically connected architecture is given in Fig. 9. It is important to highlight that the admittance-switch network can be regarded as a general framework, which enables single/fully/group/tree/forest/band/stem-connected architectures by activating the corresponding interconnection links between ports. For example, the fully-connected architecture is realized when all switches are turned ON, and the single-connected architecture is realized when all switches are turned OFF. Beyond that, this admittance-switch network can also support various grouping strategies, different from the group/forest-connected architectures where the ports are uniformly and adjacently grouped. For example, one can control the switches to realize group-connected architectures with various group size of each group [43], or to realize group-connected architectures where each group contains interlaced elements [44], as illustrated in Fig. 10.
@@
III-D Architecture Design: Non-Reciprocal Architectures
In addition to the reciprocal architectures [26, 40, 43, 44, 41] summarized in the previous subsection, there are also non-reciprocal architectures of BD-RIS realized by non-reciprocal circuits, e.g., isolators, gyrators, circulators [45, 46, 47]. Specifically, [47] proposes a physics-consistent device model, where all the $M$ elements are uniformly divided into $G$ groups with each containing $\bar{M}$ elements interconnecting with a non-reciprocal device. In each group, each element is modeled as a 2-port reciprocal network where one port interacts with free space and the other is connected to a non-reciprocal $\bar{M}$-port circuit. Two illustrative examples for a 2-element interconnected architecture realized by the isolator/gyrator and a 3-element interconnected architecture realized by the circulator are respectively given in Fig. 11(a) and Fig. 11(b). Due to the introduction of non-reciprocal circuits in each interconnected group, the resulting scattering matrix for a lossless network is a block-diagonal matrix, i.e., $\mathbf{\Theta}=\mathsf{blkdiag}(\mathbf{\Theta}_{1},\ldots,\mathbf{\Theta}_{G})$ with each block being asymmetric and unitary, that is
@@ III-D Architecture Design: Non-Reciprocal Architectures
One extreme case corresponds to the most flexible non-reciprocal BD-RIS whose scattering matrix is simply a full and unitary matrix. Another extreme case corresponds to the non-diagonal RIS [45], whose scattering matrix is asymmetric and contains only $M$ nonzero entries. In this case, $\mathbf{\Theta}$ has the following form
@@
where, for instance, $[\mathbf{\Theta}]_{1,2}\neq 0$ controls the signal flow from element 2 to element 1, that is, the signal impinging on element 2 is purely reflected by element 1.
III-E Unified Modes and Architectures
@@
Example 1: Hybrid Mode Realized by Group-Connected Architecture. In this case, every two elements are back to back placed and interconnected with each other via a 2-port fully-connected architecture to construct a cell, and every $\frac{\bar{M}}{2}$ cells are interconnected with each other to construct a group, as illustrated in Figs. 12(a) and 13(a). Therefore, we have the scattering matrix $\mathbf{\Theta}$ blocked as
@@
with $\mathbf{\Theta}_{i,j,g}\in\mathbb{C}^{\frac{\bar{M}}{2}\times\frac{\bar{M}}{2}}$, $\forall g\in\mathcal{G}$. When there is only one source located within the coverage of sector 1 as illustrated in Fig. 12(a), we introduce more intuitive notations, i.e., $\mathbf{\Theta}_{\mathsf{r}}=\mathbf{\Theta}_{1,1}$, $\mathbf{\Theta}_{\mathsf{t}}=\mathbf{\Theta}_{2,1}$, $\mathbf{\Theta}_{\mathsf{r},g}=\mathbf{\Theta}_{1,1,g}$ and $\mathbf{\Theta}_{\mathsf{t},g}=\mathbf{\Theta}_{2,1,g}$, $\forall g\in\mathcal{G}$ to describe the power reflecting (from sector 1 to sector 1) and transmitting (from sector 1 to sector 2) [36]. Then, when the group-connected reconfigurable impedance network is lossless, i.e., $\mathbf{\Theta}^{\mathsf{H}}\mathbf{\Theta}=\mathbf{I}_{M}$ and $\mathbf{\Theta}=\mathbf{\Theta}^{\mathsf{T}}$, we have the following constraint:
@@
which means the sum of the reflected power and transmitted power is conserved without loss. Specifically when $G=\frac{M}{2}$, both $\mathbf{\Theta}_{\mathsf{r}}$ and $\mathbf{\Theta}_{\mathsf{t}}$ are diagonal matrices, which correspond to the STARS, STAR-RIS or IOS [37].
@@
Example 2: Multi-Sector Mode Realized by Group-Connected Architecture. In this case, every $L$ elements are located at the edge of a polygon and interconnected with each other via an $L$-port fully-connected architecture to construct a cell, as illustrated in Figs. 12(b) and 13(b). Here we focus on the case where cells are isolated from each other for simplicity, while we can also have inter-cell interconnections as in the hybrid mode. Therefore, we have the scattering matrix $\mathbf{\Theta}$ blocked as
@@
with $\varTheta_{i,j,n}\in\mathbb{C}$, $\forall n=1,\ldots,\frac{M}{L}$. When there is only one source located within the coverage of sector 1, as illustrated in Fig. 12(b), we focus only on the first block-column of matrix $\mathbf{\Theta}$, i.e., $\mathbf{\Theta}_{l,1}$, $\forall l\in\mathcal{L}$, which describes the power scattering from sector 1 to all other sectors [38]. Then, when the reconfigurable impedance network is lossless, i.e., $\mathbf{\Theta}^{\mathsf{H}}\mathbf{\Theta}=\mathbf{I}_{M}$, we have the constraint:
@@
A special case with $L=2$ corresponds to the STARS, STAR-RIS, or IOS [37].

### v2

@@ III-D.2 Fully-Connected
Figure 9: Examples of reciprocal architectures: (a) A 4-element D-RIS with single-connected architectures; (b) a 4-element BD-RIS with fully-connected architectures; (c) a 6-element BD-RIS with group-connected architectures (each group containing 3 interconnected elements); (d) a 4-element tridiagonal BD-RIS with tree-connected architectures; (e) a 5-element arrowhead BD-RIS with tree-connected architectures; (f) a 6-element BD-RIS with forest-connected architectures (each group containing 3 interconnected elements).
III-D.3 Group-Connected
The circuit complexity (characterized by the required number of reconfigurable admittance components) of the fully-connected architecture grows quadratically with the number of elements $M$. This makes it difficult to implement fully-connected BD-RISs in practice especially when a large dimension is required. To trade the circuit complexity and wave manipulation flexibility of RIS, a group-connected architecture has been proposed in [shen2021]. In this architecture, the $M$ ports in the reconfigurable impedance network are uniformly divided into $G$ groups, and every $\bar{M}=\frac{M}{G}$ ports within one group are connected to each other to construct a fully-connected architecture. Specifically, in group $g$, $\forall g\in\mathcal{G}=\{1,\ldots,G\}$, each port $m_{g}=(g-1)\bar{M}+m$ is connected to ground via it own reconfigurable admittance component $Y_{m_{g}}$ and to another port $^{\prime}_{g}=(g-1)\bar{M}+m^{\prime}, mm$$^{\prime}\neq m$, $\forall m,m^{\prime}\in\bar{\mathcal{M}}=\{1,\ldots,\bar{M}\}$, via $Y_{m_{g},m^{\prime}_{g}}$, which satisfies $Y_{m_{g},m^{\prime}_{g}}=Y_{m^{\prime}_{g},m_{g}}$ for $m^{\prime}>m$. This indicates that the group-connected architecture is realized by in total $G\frac{\bar{M}(\bar{M}+1)}{2}=\frac{M(\bar{M}+1)}{2}$ reconfigurable admittance components. An illustrative example for a 6-element BD-RIS with group-connected architecture is given in Fig. 9(c). Hence, $\mathbf{Y}_{I}$ is a block-diagonal matrix
@@ III-D.3 Group-Connected
where each $\mathbf{Y}_{I,g}\in\mathbb{C}^{\bar{M}\times\bar{M}}$ is a full and symmetric matrix. Accordingly, $\mathbf{Z}_{I}$ is a block-diagonal matrix with each block being a full and symmetric matrix. By (16), for a lossless network, its scattering matrix $\mathbf{\Theta}$ is a block-diagonal matrix, with each block, $\mathbf{\Theta}_{g}\in\mathbb{C}^{\bar{M}\times\bar{M}}$, being symmetric and unitary, i.e.,
@@
III-D.4 Tree-Connected
While fully-connected architecture has the highest wave manipulation flexibility to achieve optimal performance, this is achieved at the cost of complicated circuit topology design requiring numerous reconfigurable admittance components. To reduce the circuit complexity while maintaining the optimal (i.e., same as fully-connected) performance in single-user multiple input single output (MISO) systems, tree-connected architecture where the circuit topology forms a tree based on graph theory [bondy2008graph] has been proposed in [nerini2024beyond]. In this architecture, the circuit complexity reduces to in total 2M12M-1 admittance components, which can be much less than that of the fully-connected architecture. Two typical tree-connected architectures, leading to respectively tridiagonal and arrowhead admittance matrices, have been introduced in [nerini2024beyond] and are reviewed as follows.
  • Tridiagonal: In this case, each port $m$ is connected to ground via its own admittance $Y_{m}$ and to port $m+1$ via an admittance $Y_{m,m+1}$, $\forall m\in\mathcal{M}\backslash\{M\}$. An illustrative example for a 4-element BD-RIS with such architecture is given in Fig. 9(d). This mathematically leads to a tridiagonal admittance matrix
@@ III-D.4 Tree-Connected
III-D.5 Forest-Connected
Similar to the extension from fully-connected to group-connected architectures, the circuit complexity of the tree-connected architecture can be further reduced by dividing the $M$ ports into $G$ groups and constructing each group as a $\bar{M}$-port tree-connected architecture. This is referred to as the forest-connected architecture [nerini2024beyond] with a circuit complexity $G(2\bar{M}-1)$. An illustrative example for a 6-element RIS with forest-connected architecture is given in Fig. 9(f). Hence, $\mathbf{Y}_{I}$ is a block-diagonal matrix with each block being a symmetric and tridiagonal (or arrowhead) matrix. Two extreme cases with $G=M$ and $G=1$ respectively correspond to single-connected and tree-connected architectures. For a lossless network with the forest-connected architecture, its impedance matrix $\mathbf{Z}_{I}$ also has a block-diagonal shape with each block being full and symmetric, and its scattering matrix $\mathbf{\Theta}$ is generally a block-diagonal matrix with each block being symmetric and unitary. Therefore, using $\mathbf{\Theta}$ or $\mathbf{Z}_{I}$ is, gain, not sufficient to fully reflect the mathematical constraint of the forest-connected architecture.
III-D.6 Band- and Stem-Connected
Inspired by the tree-connected architecture that can perfectly match the performance of fully-connected architecture in single-user MISO systems with significantly reduced circuit complexity, the optimal architectures that can perfectly match the performance of fully-connected architecture in multi-user MIMO systems have been proposed in [wu2025beyond]. In the optimal architectures, each port is connected to ground via its own reconfigurable admittance component, while the number of interconnections between ports is theoretically constrained by the DoF in multi-user MIMO systems, and is generally much less than that of fully-connected architectures. Two representative examples which could construct optimal architectures, namely band-connected [wu2025beyond] and stem-connected [wu2025beyond, zhou2024novel], are reviewed as follows.
  • Band-Connected: In this case, each port $m$ is connected to ground via $Y_{m}$ and to the following $q$ ports, i.e., $m+1,\ldots,m+q$ (if any), via corresponding admittance components, where $q$ denotes the band width. An illustrative example for a 4-element BD-RIS with band-connected architecture ($q=2$) is given in Fig. 10(a). This mathematically leads to a band admittance matrix
@@ III-D.6 Band- and Stem-Connected
    Figure 10: Examples of band- and stem-connected architectures: (a) a 4-element BD-RIS with band-connected architectures (band width $q=2$); (b) a 5-element BD-RIS with stem-connected architectures (stem width $q=2$).
  • Stem-Connected: In this case, each port is connected to ground via $Y_{m}$. Meanwhile, there are $q$ ports with $q$ denoting the stem width, marked as $c_{i},\forall i\in\{1,\ldots,q\}$, each of which connects to all other ports via corresponding admittance components. An illustrative example for a 5-element BD-RIS with such architecture ($q=2$, $c_{1}=1$, $c_{2}=2$) is given in Fig. 10(b). Assuming $c_{i}\in\{1,\ldots,q\}$, we have an admittance matrix with the following structure:
@@
It is worth noting that, similar to tree- and forest-connected architectures, for a lossless network with band- and stem-connected architectures, its impedance matrix $\mathbf{Z}_{I}$ and scattering matrices $\mathbf{\Theta}$ will both be full matrices [^3] Note that all the reciprocal architectures we talk about are explicitly described by the admittance parameter, instead of the scattering parameter. For example, the band-connected architecture has a band admittance matrix, where each nonzero entry in the off-diagonal indicates two ports are connected. However, the resulting scattering matrix will not be a band matrix, but instead a full matrix. This is also physically correct since a band-connected architecture essentially makes all the ports connected to each other, such that all entries in the scattering matrix can contribute to the wave scattering.. Therefore, $\mathbf{Y}_{I}$ should be used to accurately reflect the constraint of these two architectures.
III-D.7 Dynamically Connected
@@
In addition to the aforementioned fixed architectures which do not adapt to the channel environment (i.e., the admittance components are tuned but the circuit topology dictated by the interconnections is fixed), a dynamically connected architecture whose circuit topology can be changed during the transmission has also been proposed in [li2023dynamic]. This can be implemented by an admittance-switch network which enables a joint control of the values of admittance components and the ON/OFF states of switches. In this architecture, every port $m$ is connected to ground via its own admittance $Y_{m}$ and connected to port $^{\prime}, m>m$$^{\prime}>m$, $\forall m,m^{\prime}\in\mathcal{M}$ by the series of an admittance $Y_{m,m^{\prime}}$ and a switch $\mathsf{SW}_{m,m^{\prime}}$. This indicates that the dynamically connected architecture is realized by in total $\frac{M(M+1)}{2}$ reconfigurable admittance components and $\frac{M(M-1)}{2}$ switches. An illustrative example for a 6-element RIS with dynamically connected architecture is given in Fig. 11. It is important to highlight that the admittance-switch network can be regarded as a general framework, which enables single/fully/group/tree/forest/band/stem-connected architectures by activating the corresponding interconnection links between ports. For example, the fully-connected architecture is realized when all switches are turned ON, and the single-connected architecture is realized when all switches are turned OFF. Beyond that, this admittance-switch network can also support various grouping strategies, different from the group/forest-connected architectures where the ports are uniformly and adjacently grouped. For example, one can control the switches to realize group-connected architectures with various group size of each group [li2023dynamic], or to realize group-connected architectures where each group contains interlaced elements [nerini2024static], as illustrated in Fig. 12.
@@
III-E Architecture Design: Non-Reciprocal Architectures
In addition to the reciprocal architectures [shen2021, nerini2024beyond, li2023dynamic, nerini2024static, wu2025beyond] summarized in the previous subsection, there are also non-reciprocal architectures of BD-RIS realized by non-reciprocal circuits, e.g., isolators, gyrators, circulators [li2022reconfigurable, li2024coordinated, xu2024non]. Specifically, [xu2024non] proposes a physics-consistent device model, where all the $M$ elements are uniformly divided into $G$ groups with each containing $\bar{M}$ elements interconnecting with a non-reciprocal device. In each group, each element is modeled as a 2-port reciprocal network where one port interacts with free space and the other is connected to a non-reciprocal $\bar{M}$-port circuit. Two illustrative examples [^4] Although here we only give examples for 2-element and 3-element interconnected non-reciprical architectures for the purpose of easy illustration and simple hardware implementation, we would like to clarify that large-scale non-reciprocal architectures are practically feasible by using the combinations of multiple isolators and circulators. for a 2-element interconnected architecture realized by the isolator/gyrator and a 3-element interconnected architecture realized by the circulator are respectively given in Fig. 13(a) and Fig. 13(b). Due to the introduction of non-reciprocal circuits in each interconnected group, the resulting scattering matrix for a lossless network is a block-diagonal matrix, i.e., $\mathbf{\Theta}=\mathsf{blkdiag}(\mathbf{\Theta}_{1},\ldots,\mathbf{\Theta}_{G})$ with each block being asymmetric and unitary, that is
@@ III-E Architecture Design: Non-Reciprocal Architectures
One extreme case corresponds to the most flexible non-reciprocal BD-RIS whose scattering matrix is simply a full and unitary matrix. Another extreme case corresponds to the non-diagonal RIS [li2022reconfigurable], whose scattering matrix is asymmetric and contains only $M$ nonzero entries. In this case, $\mathbf{\Theta}$ has the following form
@@
where, for instance, $\mathbf{\Theta}]_{1,2}\neq 0$ controls the signal flow from element 2 to element 1, that is, the signal impinging on element 2 is purely reflected by element 1.
Remark 6: Based on the above illustration, we notice that both reciprocal and non-reciprocal architectures have their pros and cons. On the one hand, reciprocal architectures have advantages over non-reciprocal ones from the following two perspectives: 1) The reciprocal architectures have mathematical constraints whose locations of nonzero entries are directly reflected in the circuit topology, and this can potentially facilitate the development of various optimization methods; 2) the reciprocal architectures have simpler circuit designs without embedding non-reciprocal devices and can be more cost-effective in practical implementations. On the other hand, non-reciprocal architectures have the advantage over reciprocal ones in providing more wave manipulation flexibility due to the relaxation of the symmetric constraint in the impedance/admittance/scattering matrix of BD-RIS.
III-F Unified Modes and Architectures
@@
Example 1: Hybrid Mode Realized by Group-Connected Architecture. In this case, every two elements are back to back placed and interconnected with each other via a 2-port fully-connected architecture to construct a cell, and every $\frac{\bar{M}}{2}$ cells are interconnected with each other to construct a group, as illustrated in Figs. 14(a) and 15(a). Therefore, we have the scattering matrix $\mathbf{\Theta}$ blocked as
@@
with $\mathbf{\Theta}_{i,j,g}\in\mathbb{C}^{\frac{\bar{M}}{2}\times\frac{\bar{M}}{2}}$, $\forall g\in\mathcal{G}$. When there is only one source located within the coverage of sector 1 as illustrated in Fig. 14(a), we introduce more intuitive notations, i.e., $\mathbf{\Theta}_{\mathsf{r}}=\mathbf{\Theta}_{1,1}$, $\mathbf{\Theta}_{\mathsf{t}}=\mathbf{\Theta}_{2,1}$, $\mathbf{\Theta}_{\mathsf{r},g}=\mathbf{\Theta}_{1,1,g}$ and $\mathbf{\Theta}_{\mathsf{t},g}=\mathbf{\Theta}_{2,1,g}$, $\forall g\in\mathcal{G}$ to describe the power reflecting (from sector 1 to sector 1) and transmitting (from sector 1 to sector 2) [li2022beyond]. Then, when the group-connected reconfigurable impedance network is lossless, i.e., $\mathbf{\Theta}^{\mathsf{H}}\mathbf{\Theta}=\mathbf{I}_{M}$ and $\mathbf{\Theta}=\mathbf{\Theta}^{\mathsf{T}}$, we have the following constraint:
@@
which means the sum of the reflected power and transmitted power is conserved without loss. Specifically when $G=\frac{M}{2}$, both $\mathbf{\Theta}_{\mathsf{r}}$ and $\mathbf{\Theta}_{\mathsf{t}}$ are diagonal matrices, which correspond to the STARS, STAR-RIS or IOS [zhang2022intelligent].
@@
Example 2: Multi-Sector Mode Realized by Group-Connected Architecture. In this case, every $L$ elements are located at the edge of a polygon and interconnected with each other via an $L$-port fully-connected architecture to construct a cell, as illustrated in Figs. 14(b) and 15(b). Here we focus on the case where cells are isolated from each other for simplicity, while we can also have inter-cell interconnections as in the hybrid mode. Therefore, we have the scattering matrix $\mathbf{\Theta}$ blocked as
@@
with $\varTheta_{i,j,n}\in\mathbb{C}$, $\forall n=1,\ldots,\frac{M}{L}$. When there is only one source located within the coverage of sector 1, as illustrated in Fig. 14(b), we focus only on the first block-column of matrix $\mathbf{\Theta}$, i.e., $\mathbf{\Theta}_{l,1}$, $\forall l\in\mathcal{L}$, which describes the power scattering from sector 1 to all other sectors [li2023beyond]. Then, when the reconfigurable impedance network is lossless, i.e., $\mathbf{\Theta}^{\mathsf{H}}\mathbf{\Theta}=\mathbf{I}_{M}$, we have the constraint:
@@
A special case with $L=2$ corresponds to the STARS, STAR-RIS, or IOS [zhang2022intelligent].